OpenStudy (anonymous):

which of the following has both 1 and -3 as the solution? A. x2 − 2x − 3 = 0 B. x2 + 2x − 3 = 0 C. x2 − 4x + 3 = 0 D. x2 + 4x + 3 = 0 math has always been my weaker subject in high school. anyone please help. college is hurting me with math

OpenStudy (ranga):

Just substitute x = 1 in each equation and see if it satisfies that equation. Then try -3. The equation that satisfies both is the answer.

OpenStudy (anonymous):

thank you. math has always been my weaker subject

OpenStudy (ranga):

For example, the first equation is: x^2 − 2x − 3 = 0 Put x = 1 1^2 - 2(1) - 3 = 1 - 2 - 3 = -4 (It is not 0. Therefore x = 1 does NOT satisfy this equation. Move on to the next. First put x = 1 and calculate the left hand side of the equation. If it is zero them try x = -3 in the same equation. If the same equation makes the left hand side go to 0 when x = 1 AND x = -3 you would have found the answer.

OpenStudy (anonymous):

okay y=2(x+3)(x-4) what is the y intercept? this is all so confusing its been almost 6 years since high school. idk how to figure this out i need explanation

OpenStudy (anonymous):

would you put whatever a b c d are into y? and try and solve?

OpenStudy (ranga):

y-intercept means where the graph will cut the y-axis. On the y axis, x = 0. Therefore, to find the y-intercept, put x = 0 and find y. That will be your y-intercept. y=2(x+3)(x-4) To find y-intercept, put x = 0 y = 2(0+3)(0-4) = 2 * 3 * (-4) = -24

OpenStudy (anonymous):

wow you make it seem so easy. how do you know so much?